# Matroid variant of Matiyasevich formula and its application

### v1

@@ Abstract
In 1977, Yu.V. Matiyasevich proposed a formula expressing the chromatic polynomial of an arbitrary graph as a linear combination of flow polynomials of subgraphs of the original graph. In this paper, we prove that this representation is a particular case of one (easily verifiable) formula, namely, the representation of the characteristic polynomial of an arbitrary matroid as a linear combination of characteristic polynomials of dual matroids. As applications, we consider an explicit expression for the flow polynomial of a complete graph and a formula for the characteristic polynomial of the matroid dual to the matroid of the projective geometry over a finite field. We prove, in particular, that their major coefficients are defined by the beginning of a certain row in the Pascal triangle. We study in detail the connection with convolution formulas and other results for Tutte polynomials.
@@
Keywords: dual matroid, characteristic polynomial, flow polynomial, Matiyasevich formula, $PG(n,q)$, Tutte polynomial, convolution formula.
@@ 1 The structure of the paper, the main theorem
The main goal of this paper is to provide a very simple derivation of the formula that expresses the characteristic polynomial of an arbitrary matroid as a linear combination of characteristic polynomials of dual matroids and to consider applications of this formula. As particular cases, this representation includes the Matiyasevich formula [12], which expresses the flow polynomial of an arbitrary graph as a combination of chromatic polynomials of its subgraphs, as well as its dual formula, which uses factor graphs. We can treat both of these formulas as corollaries of one well-known correlation between vacuum Feynman amplitudes over a residue ring [9][8]. With the help of Feynman amplitudes one can also represent a flow polynomial as the sum of Legendre symbols [8]. However, in this case, everything is much simpler. For deriving a generalized variant of the Matiyasevich formula in a more convenient matroid language, we need only elementary combinatorics.
@@
An analogous situation occurs in calculating the characteristic polynomial of the matroid which is dual to $PG(n,q)$, i. e., the matroid of an $n$-dimensional projective space over a finite field $\mathbb{F}_{q}$. The characteristic polynomial of the matroid $PG(n,q)$ has roots $1,q,q^{2},\ldots,q^{n-1}$ and therefore the expression for it is evident, as distinct from the characteristic polynomial of the dual matroid. In factorization of the matroid $PG(n,q)$ in a nontrivial case there occur matroids $PG(k,q)$, $k\leq n$; their number equals the $q$-binomial coefficient (${n\choose k}_{q}$. As a result, we obtain an explicit expression for the characteristic polynomial of the matroid dual to the matroid $PG(n,q)$.
We describe these results in the next section, together with the Matiyasevich formula (deduced as a corollary of the main theorem) and the main results obtained in [11]. We observe that major coefficients of all characteristic polynomials of matroids considered in Section 2 (matroids dual to matroids of projective geometries, uniform matroids, bond matroids of complete graphs, i. e., their flow polynomials) coincide (accurate to alternating signs) with beginnings of rows of the Pascal triangle. In Section 3, we study the connection of the main theorem with convolution formulas for Tutte polynomials obtained earlier in papers [5] and [7]. In this section, we tried to thoroughly systemize results obtained for Tutte polynomials in many papers; their more concrete statements with certain additions have allowed us to obtain formulas given in sections 1 and 2. The author of this paper is most impressed by the paper [18] published by Tutte in 1967. As appeared, the formulas for Tutte polynomials of complete graphs given in the mentioned paper are the same as correlations for flow polynomials obtained by us in Section 2 of this paper. Moreover, Tutte considers the exponential generating function for Tutte polynomials of complete graphs. This function allows one to obtain a short expression for the flow polynomial of a complete graph in terms of the differentiation operation.
@@
Replacing $A$ with $E-B$ and doing the matroid $M$ in formula (8) with $M^{*}$ (recall that $\left(M^{*}\right)^{*}=M$) we get formula (7). Therefore, it suffices to prove only correlation (6). Performing substitution (2) and putting $A=E-B$ in the left-hand side of formula (6), we turn it into the form
@@
$\sum_{A\subseteq E}q^{-r(A)}\zeta_{q}(1)^{|A|}\sum_{B:A\subseteq A\subseteq E}\zeta_{q}(1)^{|B|-|A|}(-1)^{|E|-|B|}=$ (10)
$=$ $\sum_{A\subseteq E}q^{-r(A)}\zeta_{q}(1)^{|A|}(\zeta_{q}(1)-1)^{|E|-|A|}.$
@@ 2 Applications of Theorem 1 and Corollary 1
The characteristic polynomial of the dual matroid $M^{*}_{G}$ is the so-called flow polynomial $F_{G}$ [17]. Let us fix some orientation of edges in $E$ and denote by the symbol $\mathbb{Z}_{q}$ the additive group of integers modulo $q$. In what follows, we understand a nonzero flow as a correspondence between each directed edge $e\in E$ and a nonzero element $k_{e}$ in $\mathbb{Z}_{q}$ such that for each vertex $v\in V$ the sum $\sum k_{e}$ for all $e$ that enter in $v$ coincides with the analogous sum for all $e$ outgoing $v$. The number of everywhere nonzero flows is independent of the orientation of edges; it equals the value of the flow polynomial $F_{G}(q)$.
@@
Let us apply formula (12) for $G=K_{n}$. All terms in the right-hand side for various representations of $V$ as the union of disjoint parts $V_{1},V_{2},\ldots,V_{m}$ are the same, provided that their powers form one and the same partition $\lambda$ of the number $n$. By Lemma 1 the quantity of such representations equals $f(\lambda)$. The number of edges in the subgraph $H$ such that each their connectivity component represents a complete subgraph with $\lambda_{i}$ vertices obeys the evident formula
@@
Let $N=n(n-1)/2$ and $N^{\prime}=(n-1)(n-2)/2$. The major $n-2$ powers of the flow polynomial $F_{K_{n}}(x)$ take the form
@@
As an illustration, let us consider the result of the calculation of the flow polynomial of a complete graph with 5 vertices
@@
Let us now discuss the application of the main theorem in cases when the matroid $M$ is not, generally speaking, a graphic one. Recall that [15] the uniform matroid $U_{m,n}$, $m\leq n$, $m,n\in{\mathbb{N}}_{0}$, is defined on the ground set $E$, $|E|=n$, by the rank function $r$ as follows:
@@
Let us thoroughly study the application of formula (11) in the case of finite projective geometries, i. e., projective spaces over a finite field ${\mathbb{F}}_{q}$. Recall that elements of the ground set $E$ of the matroid $PG(n-1,q)$ are all distinct points of the projective space of dimension $n-1$ over the field ${\mathbb{F}}_{q}$ (see details in [15, Ch. 6]). The rank of the subset $A$ of the set $E$ is the dimension of the linear space spanned by $A$. One can easily make sure that $|E|=\frac{q^{n}-1}{q-1}={n\choose 1}_{q}$; here the Gaussian binomial coefficient (${n\choose k}_{q}$, $n,k\in\mathbb{N}_{0}$, obeys the formula
@@
The following assertion is a generalization of the remark about the number of elements in the ground set ([15, Proposition 6.1.4]):
@@
In addition, it is well known that (see [21, Proposition 7.5.3] and references to fundamental works by Richard Stanley mentioned there) the characteristic polynomial of the matroid $PG(n-1,q)$ takes the form
@@ 3 The Tutte polynomial and the matroid variant of the Matiyasevich formula
See fundamentals of the theory of Tutte polynomials, in particular, in [20], [3], [21], and in the book [19] by Tutte (in the case of graphic matroids). Specific properties of characteristic polynomials that are of interest for us are described in classical papers [17] and [21]. The review [13] has appeared to be very useful for us as a reference book of formulas for various Tutte polynomials.
@@
$=$ $(x-1)^{r(E)}\left(1+\frac{1}{(x-1)^{|E|}}\right)=x^{|E|}(x-1)^{r(E)-|E|}.$
@@
Therefore, with $0<|A|\leq m$ we get the equality $T_{U_{m,n}|A}(0,y)=0$, and with $m\leq|A|<n$ we conclude that $T_{U_{m,n}/A}(x,0)=0$. Therefore, among all terms in the sum in the right-hand side of the equality, only those ones that correspond to cases $A=\emptyset$ and $A=E$ differ from zero. ∎
@@
In the paper [16], V. Reiner generalizes equality (21) to a certain two-parameter family of equalities, where formula (21) is a limit correlation. In  [6], J. P. A. Kung establishes some two-parameter identities for characteristic polynomials of matroids. Finally, in [7] J. P. S. Kung obtains several identities which generalize identities established by him earlier and the convolution formula (21). Identity 3 in this paper takes the form
@@
Another way to derive the same expression (25) together with the exponential generating function for Tutte polynomials of projective geometries over the field $\mathbb{F}_{q}$ is proposed in the paper [2]. Evidently, the value of the $n$th derivative of the exponential generating function at zero coincides with the $n$th term of the series. Below we apply such an approach for deriving an explicit formula for the characteristic polynomial, calculating the flow polynomial of the complete graph in a different way. The technique used by us is based on results obtained by Tutte in [18].
@@
here $c(G)$ is the number of connectivity components of the graph $G$. For the dichromatic polynomial of a complete graph, Tutte has obtained an equation ([18, equation (16)]), which in terms used in Section 2 takes the form
@@
The previous correlation was strengthened by Tutte as the following formula for the exponential generating function for $Q_{K_{n}}(u,v)v^{n}$ ([18, equation (17)]):
@@
Recall that the set $A$: $A\subseteq E$ of the matroid $M=(E,r)$ is said to be independent, if $r(A)=|A|$. If, in addition, $r(A)=r(E)$, then the independent set is called a base. The minimal set, which is not independent, is called a cycle of the matroid $M$. A cycle of the matroid $M^{*}$ is called a cocycle. Let us associate all elements $e$, $e\in E$, with distinct positive integers $l_{e}$. A cycle $C$ without an element $e^{\prime}$ such that $l_{e^{\prime}}=\max_{e\in C}l_{e}$ is called a broken circuit. In [17], G. C. Rota has obtained the following result ([17, P. 359, Corollary 1]):
@@ 4 Conclusion
In this paper, we propose a matroid form of one variant of the Matiyasevich formula and describe its application for deriving characteristic polynomials of bond matroids and matroids dual to matroids of finite projective geometries. One can deduce most results from correlations for the corresponding Tutte polynomials. However, the use of one variable in the case, when two of them are not necessary, seems to give an easier way to obtain such formulas. On the other hand, this way allows us to better understand (in a sense) the interconnection of two variables in the Tutte polynomial. It also allows us to clarify the structure of major powers of considered polynomials.

### v2

@@ Abstract
In 1977, Yu. V. Matiyasevich proposed a formula expressing the chromatic polynomial of an arbitrary graph as a linear combination of flow polynomials of subgraphs of the original graph. In this paper, we prove that this representation is a particular case of one (easily verifiable) formula, namely, the representation of the characteristic polynomial of an arbitrary matroid as a linear combination of characteristic polynomials of dual matroids. As an application, we represent the flow polynomial of a complete graph with $n$ vertices as the sum of elementary terms with respect to all partitions of positive integer $n$. Since the growth rate of the number of all partitions is less than exponential, this technique allows us to evaluate the flow polynomial for values of $n\approx 50$. We also get an explicit expression for the characteristic polynomial of the matroid dual to the matroid of the projective geometry over a finite field. We prove, in particular, that major coefficients of all these polynomials coincide with the beginning of the row in the Pascal triangle, whose number equals the quantity of elements in the corresponding matroid. At the end part of the paper, we consider one more approach, which allows us to obtain the same results of application of our main theoren by using properties of the Tutte polynomial and the classical Rota formula for coefficients of the characteristic polynomial of a matroid. In addition, we describe the connection between the matroid variant of the Matiyasevich formula and convolution formulas for Tutte polynomials.
@@
Keywords: dual matroid, characteristic polynomial, chromatic and flow polynomial, Matiyasevich formula, $PG(n,q)$, Tutte polynomial, convolution formula.
@@ 1 The structure of the paper, the main theorem
1.1 Motivation and the paper structure
The main goal of this paper is to provide a very simple derivation of the formula that expresses the characteristic polynomial of an arbitrary matroid as a linear combination of characteristic polynomials of dual matroids and to consider applications of this formula. As particular cases, this representation includes the Matiyasevich formula [11], which expresses the flow polynomial of an arbitrary graph as a combination of chromatic polynomials of its subgraphs, as well as its dual formula, which uses factor graphs. We can treat both of these formulas as corollaries of one well-known correlation between vacuum Feynman amplitudes over a residue ring [8, 10]. With the help of Feynman amplitudes one can also represent a flow polynomial as the sum of Legendre symbols [7]. However, in this case, everything is much simpler. For deriving a generalized variant of the Matiyasevich formula in a more convenient matroid language, we need only elementary combinatorics.
@@ 1.1 Motivation and the paper structure
An analogous situation occurs in calculating the characteristic polynomial of the matroid which is dual to $PG(n-1,q)$, i. e., the matroid of an $(n-1)$-dimensional projective space over a finite field $\mathbb{F}_{q}$. The characteristic polynomial of the matroid $PG(n-1,q)$ has roots $1,q,q^{2},\ldots,q^{n-1}$ and therefore the expression for it is evident, as distinct from the characteristic polynomial of the dual matroid. In factorization of the matroid $PG(n-1,q)$ in a nontrivial case there occur matroids $PG(k-1,q)$, $k\leq n$; their number equals the $q$-binomial coefficient (${n\choose k}_{q}$. As a result, we obtain an explicit expression for the characteristic polynomial of the matroid dual to the matroid $PG(n-1,q)$.
We describe these results in the section 2, together with the Matiyasevich formula (deduced as a corollary of the main theorem) and the main results obtained in [10]. We observe that major coefficients of all characteristic polynomials of matroids considered in Section 2 (matroids dual to matroids of projective geometries, uniform matroids, bond matroids of complete graphs, i. e., their flow polynomials) coincide (accurate to alternating signs) with beginnings of rows of the Pascal triangle. In Section 3, we study the connection of the main theorem with convolution formulas for Tutte polynomials obtained earlier in papers [5] and [6]. In this section, we tried to thoroughly systemize results obtained for Tutte polynomials in many papers; their more concrete statements with certain additions have allowed us to obtain formulas given in sections 1 and 2. The author of this paper is most impressed by the paper [20] published by Tutte in 1967. As appeared, the formulas for Tutte polynomials of complete graphs given in the mentioned paper are the same as correlations for flow polynomials obtained by us in Section 2 of this paper. Moreover, Tutte considers the exponential generating function for Tutte polynomials of complete graphs. This function allows one to obtain a short expression for the flow polynomial of a complete graph in terms of the differentiation operation.
@@
At the end part of this paper, in Conclusion, we discuss references and further research directions.
1.2 The main theorem
@@
Replacing $A$ with $E-B$ and doing the matroid $M$ in formula (8) with $M^{*}$ (recall that $\left(M^{*}\right)^{*}=M$) we get formula (7). Therefore, it suffices to prove only correlation (6).
Performing substitution (2) and putting $A=E-B$ in the left-hand side of formula (6), we turn it into the form
@@
$\sum_{A\subseteq E}q^{-r(A)}\zeta_{q}(1)^{|A|}\sum_{B:A\subseteq B\subseteq E}\zeta_{q}(1)^{|B|-|A|}(-1)^{|E|-|B|}=$ (10)
$=$ $\sum_{A\subseteq E}q^{-r(A)}(-\zeta_{q}(1))^{|A|}(\zeta_{q}(1)-1)^{|E|-|A|}.$
@@ 2.1 Applications in the case of the graphic matroid
The characteristic polynomial of the dual (bond) matroid $M^{*}_{G}$ is the so-called flow polynomial $F_{G}$ [17]. Let us fix some orientation of edges in $E$ and denote by the symbol $\mathbb{Z}_{q}$ the additive group of integers modulo $q$. In what follows, we understand a nonzero flow as a correspondence between each directed edge $e\in E$ and a nonzero element $k_{e}$ in $\mathbb{Z}_{q}$ such that for each vertex $v\in V$ the sum $\sum k_{e}$ for all $e$ that enter in $v$ coincides with the analogous sum for all $e$ outgoing $v$. The number of everywhere nonzero flows is independent of the orientation of edges; it equals the value of the flow polynomial $F_{G}(q)$.
@@
Let us apply formula (12) for $G=K_{n}$. All terms in the right-hand side for various representations of $V$ as the union of disjoint parts $V_{1},\ldots,V_{m}$ are the same, provided that their powers form one and the same partition $\lambda$ of the number $n$. By Lemma 1 the quantity of such representations equals $f(\lambda)$. The number of edges in the subgraph $H$ such that each their connectivity component represents a complete subgraph with $\lambda_{i}$ vertices obeys the evident formula
@@
Let $n>2$, $N=n(n-1)/2$ and $N^{\prime}=(n-1)(n-2)/2$. The major $n-2$ powers of the flow polynomial $F_{K_{n}}(x)$ take the form
@@
As an illustration, let us consider the result of the calculation of the flow polynomial of a complete graph with 5 vertices:
@@
2.2 Applications in the non-graphic case
Let us now discuss the application of the main theorem in cases when the matroid $M$ is not, generally speaking, a graphic one. Recall that ([14]) the uniform matroid $U_{m,n}$, $m\leq n$, $m,n\in{\mathbb{N}}_{0}$, is defined on the ground set $E$, $|E|=n$, by the rank function $r$ as follows:
@@ 2.2 Applications in the non-graphic case
Let us thoroughly study the application of formula (11) in the case of finite projective geometries, i. e., projective spaces over a finite field ${\mathbb{F}}_{q}$. Recall that elements of the ground set $E$ of the matroid $PG(n-1,q)$ are all distinct points of the projective space of dimension $n-1$ over the field ${\mathbb{F}}_{q}$ (see details in [14, Ch. 6]). The rank of the subset $A$ of the set $E$ is the dimension of the linear space spanned by $A$, in particular $r(E)=n$. One can easily make sure that $|E|=\frac{q^{n}-1}{q-1}={n\choose 1}_{q}$; here the Gaussian binomial coefficient (${n\choose k}_{q}$, $n,k\in\mathbb{N}_{0}$, obeys the formula
@@
The following assertion is a generalization of the remark about the number of elements in the ground set ([14, Proposition 6.1.4]):
@@
In addition, it is well known that (see [24, Proposition 7.5.3] and references to fundamental works by Richard Stanley mentioned there) the characteristic polynomial of the matroid $PG(n-1,q)$ takes the form
@@ 3.1 Definition of the Tutte polynomial
See fundamentals of the theory of Tutte polynomials, in particular, in [22], [3], and in the book [21] by Tutte (in the case of graphic matroids). Specific properties of characteristic polynomials that are of interest for us are described in classical papers [17] and [24]. The review [12] has appeared to be very useful for us as a reference book of formulas for various Tutte polynomials.
@@
$=$ $(x-1)^{r(E)}\left(1+\frac{1}{(x-1)}\right)^{|E|}=x^{|E|}(x-1)^{r(E)-|E|}.$
@@
3.2 Theorem 1 as a corollary of the convolution formula for the Tutte polynomial
@@
Therefore, with $0<|A|\leq m$ we get the equality $T_{U_{m,n}|A}(0,y)=0$, and with $m\leq|A|<n$ we conclude that $T_{U_{m,n}/A}(x,0)=0$. Therefore, among all terms in the sum in the right-hand side of the equality (21), only those ones that correspond to cases $A=\emptyset$ and $A=E$ differ from zero. ∎
@@
In the paper [15], V. Reiner generalizes equality (21) to a certain two-parameter family of equalities, where formula (21) is a limit correlation. Finally, in [6] J. P. S. Kung obtains several identities which generalize identities established earlier and the convolution formula (21) in particular. Identity 3 in this paper takes the form
@@
3.3 Explicit formulas for the Tutte polynomials of some matroids and theorems of Section 2
@@
Another way to derive the same expression (25) together with the exponential generating function for Tutte polynomials of projective geometries over the field $\mathbb{F}_{q}$ is proposed in the paper [2]. Evidently, the value of the $n$th derivative of the exponential generating function at zero coincides with the $n$th term of the series. Below we apply such an approach for deriving an explicit formula for the characteristic polynomial of matroid, calculating the flow polynomial of the complete graph in a different way. The technique used by us is based on results obtained by Tutte in [20].
@@
here $c(G)$ is the number of connectivity components of the graph $G$. For the dichromatic polynomial of a complete graph, Tutte has obtained an equation ([20, equation (16)]), which in terms used in Section 2 takes the form
@@
The previous correlation was strengthened by Tutte as the following formula for the exponential generating function for $Q_{K_{n}}(u,v)v^{n}$ ([20, equation (17)]):
@@
3.4 Reason for binomiality of the major coefficients of characteristic polynomials
@@
Recall that the set $A$: $A\subseteq E$ of the matroid $M=(E,r)$ is said to be independent, if $r(A)=|A|$. If, in addition, $r(A)=r(E)$, then the independent set is called a base. The minimal set, which is not independent, is called a cycle of the matroid $M$. A cycle of the matroid $M^{*}$ is called a cocycle. Let us associate all elements $e$, $e\in E$, with distinct positive integers $l_{e}$. A cycle $C$ without an element $e^{\prime}$ such that $l_{e^{\prime}}=\max_{e\in C}l_{e}$ is called a broken circuit. G. C. Rota has obtained the following result ([17, P. 359, Corollary 1]):
@@ 4 Conclusion
In this paper, we propose a matroid form of one variant of the Matiyasevich formula and describe its application for deriving flow polynomials and characteristic polynomials of matroids dual to matroids of finite projective geometries. One can deduce most results from correlations for the corresponding Tutte polynomials. However, the use of one variable in the case, when two of them are not necessary, seems to give an easier way to obtain such formulas. On the other hand, this way allows us to better understand (in a sense) the interconnection of two variables in the Tutte polynomial. It also allows us to clarify the structure of major powers of considered polynomials.
@@
A month after the preprint of the first version of this paper was uploaded to arXiv.org, there has appeared preprint [19] by R. Takashi, where he, in particular, develops ideas of this paper and gives more references. In [16, formula (3.2)], R. C. Read and E. G. Whitehead re-derive the Matiyasevich formula. Moreover, in the same paper, they obtain an analog of this formula, which expresses a flow polynomial in terms of the sum of chromatic polynomials of factor graphs [16, formula (7.8)]. In [10], we prove the mentioned formula with the help of the Fourier transformations method (which is standard for Feynman amplitudes); in the Matiyasevich paper, this method is described regardless of Feynman amplitudes. R. C. Read and E. G. Whitehead propose a more elementary proof; previously, in Section 1.2, we describe its simplified variant for the matroid case. As applied to flow and chromatic polynomials, the indicated formula has 4 variants (we have mentioned two of them earlier; in one more variant, the flow polynomial is represented as a linear combination of chromatic polynomials of subgraphs; the fourth variant of the formula expresses the chromatic polynomial in terms of flow polynomials of factor graphs). As we have mentioned earlier, the first formula among two ones indicated in braces is formula (2.5) obtained in paper [4]. R. Takashi also gives all these formulas and their elementary proofs. Moreover, following [4] (see also [18]), R. Takashi for each formula considers its variant for a case, when each edge has a real-valued parameter, and generalizes these formulas for the matroid case. One can interpret this multiparameter variant as the partition function of multivariate Potts model .
Note that the multiparameter variants of these formulas, evidently, were obtained by D. R. Woodall in [23, Section 3] with reference to Alan Sokal. In addition, in the same paper, he also develops ideas of R. C. Read and E. G. Whitehead and gives detailed comments on them. In particular, regarding formulas (6) and (7) in the case of a graphic matroid and a bond one [23, formulas (1.1)–(1.4)] (these formulas establish a connection between the flow polynomial $F$ and the normalized chromatic polynomial denoted by the symbol $N$), D. R. Woodall concludes that “The existence of expansions of this type is not surprising, but the interchange of $N$ and $F$ is intriguing and merits further study.” One can consider this work as uncovering of this intrigue.